OpenStudy (anonymous):

So I am trying to solve a question that I cam up with, can someone please help explain it out to me? I'll fan and medal. My problem is going to be x/7 - -12=66

OpenStudy (johnweldon1993):

\[\large \frac{x}{7} - (-12) = 66\] like that?

OpenStudy (anonymous):

Yeah

OpenStudy (johnweldon1993):

Well you first need to remember that subtracting a negative turns positive..so we have \[\large \frac{x}{7} + 12 = 66\] Look better?

OpenStudy (anonymous):

Yep

OpenStudy (johnweldon1993):

So now we subtract 12 from both sides \[\large \frac{x}{7} = 54\] And now multiply both sides by 7 \[\large x = 54 \times 7\] \[\large x = ?\]

OpenStudy (anonymous):

378

OpenStudy (anonymous):

Thank you!

OpenStudy (johnweldon1993):

No problem!